OpenStudy (anonymous):

Factor: 81x^2-144x+64

OpenStudy (anonymous):

(9x-8)^2

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

It is a perfect square. Formula for a perfect square is x^2 + 2ax + a^2 = (x + a)^2

OpenStudy (anonymous):

we have a minus so our will look like this when we use the formula (9x+(-8))(9x+(-8))

OpenStudy (anonymous):

simplify it to (9x-8)^2

OpenStudy (anonymous):

So its factored completely?

OpenStudy (anonymous):

it is easier to see backwards. (9x-8)(9x-8) 81x^2-72x-72x+64 (when you foil) combine like terms (81x^2-144x+64)

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

Thanks :))

OpenStudy (anonymous):

Can you help with this one? 9-49x^2

OpenStudy (anonymous):

:) any time. You can normally tell when you are done when the degree of your polynomial is one. so just x not x^2 ... or x^#

OpenStudy (anonymous):

so (3-7x)(3+7x)

OpenStudy (anonymous):

Formula (a^2-b^2)=(a+b)(a-b)

OpenStudy (anonymous):

This only works with DIFFERENCE of squares. (when you subtract) SUM of two squares (when you add) for you is prime. So for any sum you would answer "prime" or stop factoring.

OpenStudy (anonymous):

81x^2-144x+64 (9x)^2 - 8^2 = ( 9x - 8) ( 9x + 8)